Your company makes Product 1 and Product 2 using Activity A and Activity B. The cost...

  1. Your company makes Product 1 and Product 2 using Activity A and Activity B. The cost driver for Activity A is direct labor hours and the cost driver for Activity B is number of setups. Activity A costs $125,000 and Activity B costs $675,000. If Product 1 had activity proportions of 10% for Activity A and 40% for Activity B, then
    1. what is the total indirect cost assigned to Product 1?
    2. what is the total indirect cost assigned to Product 2?
  1. Your company makes Product 1 and Product 2 using Activity A and Activity B. The cost driver for Activity A is direct labor hours and the cost driver for Activity B is number of setups. Product 1 requires 800 direct labor hours and 70 setups, but Product 2 requires 400 direct labor hours and 90 setups. If Activity A had an activity rate of $13 per direct labor hour and Activity B had an activity rate of $24 per setup, then
    1. what is the total indirect cost assigned to Product 1?
    2. what is the total indirect cost assigned to Product 2?
  1. Your company makes Product 1 using Activity A and Activity B. The cost driver for Activity A is direct labor hours and the cost driver for Activity B is number of setups. Activity A has an activity rate of $10 per direct labor hour and Activity B has an activity rate of $150 per setup. One unit of Product 1 is sold for $200, requires $30 of direct materials and $60 of direct labor, and uses 8 direct labor hours and 4 setups. If your company sells 50 units of Product 1, then what is
    1. the total manufacturing cost per unit?
    2. the gross profit per unit?
Answer #1

1.
a. Indirect costs assigned to Product 1 = $125000 x 10% + 675000 x 40% = $282,500
b. Indirect costs assigned to Product 1 = $125000 x 90% + 675000 x 60% = $517,500

2.
a. Indirect costs assigned to Product 1 = 800 x $13 + 70 x $24 = $12080
b. Indirect costs assigned to Product 1 = 400 x $13 + 90 x $24 = $7360

3.
Overhead assigned = 8 x $10 + 4 x $150 = $680
Overhead per unit = $680 / 50 = $13.60

a. Total Manufacturing Cost per unit = $30+60+13.60 = $103.60
b. Gross Profit per unit = $200 - 103.60 = $96.40 per unit

If you have any query, kindly comment with your query and please mark thumbs up.